# SV4D: Dynamic 3D Content Generation with Multi-Frame and Multi-View Consistency

### v1

@@ Abstract
We present Stable Video 4D (SV4D) — a latent video diffusion model for multi-frame and multi-view consistent dynamic 3D content generation. Unlike previous methods that rely on separately trained generative models for video generation and novel view synthesis, we design a unified diffusion model to generate novel view videos of dynamic 3D objects. Specifically, given a monocular reference video, SV4D generates novel views for each video frame that are temporally consistent. We then use the generated novel view videos to optimize an implicit 4D representation (dynamic NeRF) efficiently, without the need for cumbersome SDS-based optimization used in most prior works. To train our unified novel-view video generation model, we curated a dynamic 3D object dataset from the existing Objaverse dataset. Extensive experimental results on multiple datasets and user studies demonstrate SV4D’s state-of-the-art performance on novel-view video synthesis as well as 4D generation compared to prior works. Project page: https://sv4d.github.io.
@@ 1 Introduction
As a result, several recent techniques [48; 1; 27; 42; 68; 18; 66; 64] optimize 4D content by leveraging priors in pre-trained video and multi-view generative models via score-distillation sampling (SDS) loss [40] and its variants. However, they tend to produce unsatisfactory results due to independent modeling of object motion using video models, and novel view synthesis using multi-view generative models. In addition, they tend to take hours to generate a single 4D object due to time-consuming SDS-based optimization. A concurrent work [61] tries to partially address these issues by jointly sampling novel view videos (along both video frame and view axes) using both video and multi-view generative models; and then using the resulting novel view videos for 4D optimization. This results in a better novel view video synthesis with multi-view dynamic consistency, but several inconsistencies still remain due to the use of separate video and multi-view generative models.
In this work, we propose Stable Video 4D (SV4D) model that takes as input a single video, such as of a dynamic object, along with a user-specified camera trajectory around the object, and outputs videos of the object along each of the specified camera views. That is, given a video with $F$ number of frames and a camera trajectory with $V$ number of camera views, SV4D outputs a $V{\mkern-2.0mu\times\mkern-2.0mu}F$ grid of images, as illustrated in Fig. 1. Given a reference video, we first obtain the reference multi-views for the first frame using an off-the-shelf multi-view generator (SV3D [54]). SV4D then jointly outputs the remaining grid of images as highlighted by a yellow box in Fig. 1. In contrast to prior and concurrent works, SV4D jointly reasons along both view and motion axes, resulting in state-of-the-art multi-frame and multi-view consistency in the output novel view videos.
Specifically, we start with Stable Video Diffusion (SVD) [2], a state-of-the-art video generator, and equip it with two more attention blocks: view attention and frame attention. The view attention block aligns the multi-view images at each video frame, conditioning on the first view i.e, in the reference video. Similarly, the frame attention block aligns the multi-frame images at each view, conditioning on the first frame at each view, i.e, reference multi-view. This dual-attention design leads to significantly improved dynamic and multi-view consistency compared to prior art.
@@
Given the lack of large-scale 4D datasets, we carefully initialized SV4D weights with those from SVD [2] and SV3D [54] networks, thereby leveraging the priors learned in the existing video and multi-view diffusion models. To further train SV4D, we carefully curated a subject of Objaverse [9; 8] dataset with dynamic 3D objects, resulting in the ObjaverseDy dataset.
We perform extensive comparisons of both novel view video synthesis and 4D generation results with respective state-of-the-art methods on datasets with synthetic (ObjaverseDy, Consistent4D [18]), and real-world (DAVIS [38; 39; 4]) data. We modify the FVD metric [52] to evaluate both video frame and view consistency, validating the effectiveness and design choices of our approach. Fig. 1 shows some sample results of our approach.
@@ 2 Related Work
3D Generation. Here, we refer to the works that generate static 3D content as 3D generation. DreamFusion [40] first proposed to distill priors from the 2D diffusion model via SDS loss to optimize the 3D content from text or image. Several subsequent works [63; 50; 45; 57; 24; 59; 36; 6; 49; 43; 26; 69; 11] try to solve issues caused by the original SDS loss, such as multi-face Janus, slow generation speed, and over-saturated/smoothed generations. Recent works [16; 17; 56; 70; 58; 51] try to directly predict the 3D model of an object via a large reconstruction model. Another approach [30; 31; 32; 54; 62; 19; 23; 46; 44; 55; 29; 28] to 3D generation is generating dense multi-view images with sufficient 3D consistency. 3D content is reconstructed based on the dense multi-view images. We follow this strategy, but generate consistent multi-view videos (instead of images) and then reconstruct the 4D object.
@@
Video Generation. Recent video generation models [15; 53; 3; 2; 14; 47; 12] have shown very impressive performance with consistent geometry and realistic motions. Video generation models have good generalization capabilities, as they are trained on large-scale image and video data that are easier to collect than large-scale 3D or 4D data. Hence, they are commonly used as foundation models for various generation tasks. Well-trained video generative models have shown their potential to generate multi-view images as a 3D generator [54; 7; 13; 22; 33]. SV3D [54] adapts a latent video diffusion model (Stable Video Diffusion - SVD [2]) to generate novel multiple views. In this work, we leverage the pre-trained video generation model for 4D generation by adding an additional view attention layer to align the multiview images.
@@
4D Generation. On one hand, recent optimization-based methods [48; 1; 27; 42; 68; 18; 66; 64] can generate 4D content by distilling pre-trained diffusion models in a 4D representation [5; 20; 34] via SDS loss [40]. MAV3D [48] distills a text-to-video diffusion model and several following works [1; 27; 42; 68] combine both video and multi-view diffusion models to leverage the motion and multi-view priors. However, they tend to take hours to generate 4D content due to time-consuming optimization. On the other hand, photogrammetry-based methods [61; 37] mimic a 3D object capture pipeline by directly generating multi-frame multi-view images of a 4D content with dynamic and multi-view consistency and then directly reconstructing 4D representations with them. Although these methods have much faster speeds, inference-only pipelines are adopted due to the paucity of the 4D data, thus making the spatial-temporal consistency still unsatisfactory. In this work, we proposed a monolithic model to generate a more consistent image grid, and we use a carefully curated 4D dataset to train the model.
@@ 3 Method
Our main idea is to build multi-frame and multi-view consistency in a 4D object by surgically combining the frame-consistency in a video diffusion model, with the multi-view consistency in a multi-view diffusion model. In our case, we choose SVD [2] and SV3D [54] as the video and multi-view diffusion models respectively, for the advantages reasoned below. However, it is to be noted that any choice of attention-based diffusion models should work.
@@ 3.1 Novel View Video Synthesis via SV4D
Problem Setting. Formally, we begin with a monocular input video ${\bf J}\in\mathbb{R}^{F{\mkern-2.0mu\times\mkern-2.0mu}D}$ of a dynamic object of $F$ frames and $D\coloneq 3{\mkern-2.0mu\times\mkern-2.0mu}H{\mkern-2.0mu\times\mkern-2.0mu}W$ dimensions, where $H$ and $W$ are the height and width of each frame. Our goal is to generate an image matrix ${\bf M}\in\mathbb{R}^{V{\mkern-2.0mu\times\mkern-2.0mu}F{\mkern-2.0mu\times\mkern-2.0mu}D}$ of the 4D object consisting of $F$ dynamic frames and $V$ camera views of each frame. Similar to SV3D [54], the multi-view frames follow a camera pose trajectory $\bm{\pi}\in\mathbb{R}^{V{\mkern-2.0mu\times\mkern-2.0mu}2}=\{(e_{v},a_{v})\}_{v=1}^{V}$ as a sequence of $V$ tuples of elevation $e$ and azimuth angles $a$. We generate this image matrix by iteratively denoising samples from a learned conditional distribution $p({\bf M}|{\bf J},\bm{\pi})$, parameterized by a 4D diffusion model. Due to memory limitation in generating $V{\mkern-2.0mu\times\mkern-2.0mu}F$ images simultaneously, we break down the full generation process into multiple submatrix generation steps.
@@
SV4D Network. Our goal is to make the generated image matrix $\bf M$ dynamically consistent in the “frame” axis, and multi-view consistent in the “view” axis (see Fig. 2 left). To achieve this, we condition the image matrix generation on the corresponding frames of the monocular input video $\{{\bf M}_{0,f}\}={\bf J}$ as well as the views from reference multi-view images of the first video frame $\{{\bf M}_{v,0}\}$, which provide motion and multi-view information of the 4D object, respectively. Without loss of generality, we obtain the reference multi-view images of the first input frame by sampling from a pre-trained SV3D [54] model, which can be expressed as $p(\{{\bf M}_{v,0}\}|\{{\bf M}_{0,0}\},\bm{\pi})$. The overall SV4D novel view video synthesis can be rewritten as sampling from the distribution $p({\bf M}|\{{\bf M}_{0,f}\},\{{\bf M}_{v,0}\},\bm{\pi})$
@@
Figure 2: SV4D Model Architecture. For camera conditioning, we feed the sinusoidal embedding of camera viewpoints to the convolutional blocks in the UNet, and use the input video for cross-attention conditioning in the spatial and view attention blocks. To improve temporal consistency, we introduce an additional motion attention block, conditioned on the corresponding views of the first frame.
@@
We build the SV4D network based on SVD [2] and SV3D [54] models to combine the advantages of both video and multi-view diffusion models. As illustrated in Fig. 2, SV4D consists of a UNet with multiple layers, where each layer contains a sequence of one residual block with Conv3D layers and three transformer blocks with attention layers: spatial, view, and frame attention. Similar to SV3D, the residual Conv block takes in the noisy latents of flattened image matrix as well as handles the incorporation of conditioning camera poses $\{\bm{\pi}_{v}\}$, and the spatial attention layer handles image-level details by performing attention across the image width and height axes. To better capture motion cues in the input monocular video $\{{\bf M}_{0,f}\}$, we concatenate its VAE latents to the noisy latents ${\bf z}_{n}$ before feeding it to the UNet.
@@
To further ensure dynamic consistency across video frames, we insert a frame attention layer in each UNet block, which applies the attention mechanism in the video frame dimension. The frame attention of each novel view video is conditioned on the corresponding reference view (of the first frame) $\{{\bf M}_{v,0}\}$ via cross-attention, allowing the network to preserve dynamic coherence starting from the first frame. We initialize the weights of the frame attention layers from SVD and the rest of the network from SV3Dp, to leverage the generalizability as well as rich dynamic and multi-view priors learned from large-scale video and 3D datasets.
ObjaverseDy Dataset. Considering that there exists no large-scale training datasets with dynamic 3D objects, we curate a new 4D dataset from the existing Objeverse dataset [9; 8], a massive dataset with annotated 3D objects. Objaverse includes animated 3D objects, however, several of these animated 3D objects are not suitable for training due to having too few animated frames or insufficient motion. In addition, in the rendering stage, the common rendering and sampling setting may cause some issues. For example, dynamic objects may be out of the image if the camera distance is fixed because they have global motion; the motion of objects may be too fast or too slow if the temporal sampling step is fixed.
We follow several steps to curate and clean the 4D objects for our training purposes. We first filter out the objects based on a review of licenses. Then, we remove the objects whose animated frames are too few. To further filter out objects with minimal motion, we subsample keyframes from each video and apply simple thresholding on the maximum L1 distance between these frames as motion measurement. To render the training novel view videos, we flexibly choose the camera distance from the object. Starting from a base value, we increase the camera distance until the object fits within all frames of the rendered images. We also dynamically adjust the temporal sampling step. Starting from a base value, we increase the temporal sample step until the L1 distance between consecutive keyframes exceeds a certain threshold. These steps ensure a high-quality collection of 4D objects, with rendered multi-view videos that form our ObjaverseDy dataset.
Training Details. We train SV4D on our ObjaverseDy dataset. We choose to finetune from the SV3Dp [54] model to output 40 frames ($F=5$ frames along each of the $V=8$ views), where the parameters in the frame attention layers are initialized from SVD-xt [2]. Similar to SV3D, we train SV4D progressively by first training on the static camera orbits for 40K iterations, then fine-tuning it for 20K iterations on the dynamic orbits. We use an effective batch size of 16 during training on 2 nodes of 8 80GB H100 GPUs. For more training details, please see the appendix.
@@
Inference Sampling Scheme. Due to memory limitations, we cannot generate all the novel view frames at once when $F>5$. To generate the full $V{\mkern-2.0mu\times\mkern-2.0mu}F$ image matrix with arbitrary length videos during inference, one can naively run the submatrix generation in an auto-regressive manner. However, we observe that it often leads to severe artifacts due to error propagation during the generation process. Hence, we design a novel sampling scheme to mitigate the issue. The proposed sampling scheme is illustrated in Fig. 3. Note that we only show motion frame extensions here for illustrative purposes. We first generate a sparse set of anchor frames with SV4D (interleaved sampling) as shown in Fig. 3 (left). Then we use the anchor frames as new reference views to densely sample the remaining frames (dense sampling). To ensure a smooth transition between consecutive generations, we alternatively use the first (forward) or last (backward) anchor frame for conditioning at each diffusion step, as shown in Fig. 3 (middle). In our experiments (Fig. 8), we demonstrate that our sampling scheme can generate novel view videos that are more temporally consistent compared to using an off-the-shelf video interpolation model [25] to interpolate in between the SV4D-generated anchor frames.
Spatio-temporal CFG Scaling. During inference sampling, SVD uses a linearly increasing scale for classifier-free guidance (CFG) over time (video frame axis) to avoid losing context from the first reference frame. SV3D, on the other hand, adopts a triangular CFG scaling (in the view axis) when generating an orbital video around a static 3D object to avoid over-sharpened or over-saturated outputs. However, we observe that neither of these CFG scaling schemes suit the spatio-temporal nature in 4D image matrix generation and often cause undesirable artifacts. Hence, we propose a novel spatio-temporal CFG scaling which integrates a linearly-increasing scale in the frame axis and a triangular wave in the view axis (linearly increase when moving away from the input view, and linearly decrease when rotating back towards the input view). Details can be found in the appendix.
@@
4D Representation. As shown in Fig. 4, given the novel view videos generated by SV4D, we optimize a 4D representation to reconstruct the dynamic 3D asset. Formally, we learn a neural representation $\Psi_{\theta}:(\mathbf{x},t)\mapsto(\sigma,c)$ that maps the sampled 3D points $\mathbf{x}=(x,y,z)$ along a camera ray, and the time embedding $t$ (i.e. continuous version of “frame” in the earlier discrete case) to its volumetric density $\sigma\in\mathbb{R}_{+}$ and color $c\in\mathbb{R}^{3}_{+}$. Similar to D-NeRF [41], we represent a 4D object by the composition of a canonical NeRF that captures the static 3D appearance and a deformation field which handles the object motion across time. For each 3D point $\mathbf{x}$ in the canonical space, we trilinearly interpolate the multi-resolution hash grid features following Instant-NGP [35], and decode them as density and color via MLPσ and MLPc, respectively. The deformation field is represented by an MLP network MLPt conditioned on time embedding $t$, mapping temporal deformation of $\mathbf{x}$ to the common canonical space $\mathbf{x}+\Delta\mathbf{x}$. Overall, the dynamic NeRF parameters $\theta$ include the canonical hash grid, MLPt, MLPσ, and MLPc. Since we do not exhaustively sample temporal timestamps or dense spatial views like in prior SDS-based approaches, we observe that this dynamic NeRF representation produces better 4D results compared to other representations such as 4D Gaussian Splatting [60], which suffers from flickering artifacts and does not interpolate well across time or views.
Optimization Details. By leveraging the consistent image matrix generated by SV4D as pseudo ground-truths, we adopt a simple photometry-based optimization without the cumbersome SDS losses. The reconstruction losses include a pixel-level MSE loss, mask loss, and a perceptual LPIPS [67] loss. Similar to SV3D [54], we also use several geometric priors to regularize the output shapes, such as a mono normal loss similar to MonoSDF [65] as well as bilateral depth and normal smoothness losses [54] to encourage smooth 3D surfaces where the projected image gradients are low.

### v2

@@ Abstract
We present Stable Video 4D (SV4D) — a latent video diffusion model for multi-frame and multi-view consistent dynamic 3D content generation. Unlike previous methods that rely on separately trained generative models for video generation and novel view synthesis, we design a unified diffusion model to generate novel view videos of dynamic 3D objects. Specifically, given a monocular reference video, SV4D generates novel views for each video frame that are temporally consistent. We then use the generated novel view videos to optimize an implicit 4D representation (dynamic NeRF) efficiently, without the need for cumbersome SDS-based optimization used in most prior works. To train our unified novel view video generation model, we curate a dynamic 3D object dataset from the existing Objaverse dataset. Extensive experimental results on multiple datasets and user studies demonstrate SV4D’s state-of-the-art performance on novel-view video synthesis as well as 4D generation compared to prior works. Project page: https://sv4d.github.io.
@@ 1 Introduction
As a result, several recent techniques (Singer et al., 2023b; Bahmani et al., 2024; Ling et al., 2024; Ren et al., 2023; Zhao et al., 2023; Jiang et al., 2024c; Zeng et al., 2024; Yin et al., 2023) optimize 4D content by leveraging priors in pre-trained video and multi-view generative models via score-distillation sampling (SDS) loss (Poole et al., 2023) and its variants. However, they tend to produce unsatisfactory results due to independent modeling of object motion using video models, and novel view synthesis using multi-view generative models. In addition, they tend to take hours to generate a single 4D object due to time-consuming SDS-based optimization. Concurrent works (Yang et al., 2025; Sun et al., 2025) try to partially address these issues by jointly sampling novel view videos (along both video frame and view axes) using both video and multi-view generative models; and then using the resulting novel view videos for 4D optimization. This results in a better novel view video synthesis with multi-view dynamic consistency, but several inconsistencies still remain due to the use of separate video and multi-view generative models.
In this work, we propose Stable Video 4D (SV4D) model that takes as input a single video, such as of a dynamic object, along with a user-specified camera trajectory around the object, and outputs videos of the object along each of the specified camera views. That is, given a video with $F$ number of frames and a camera trajectory with $V$ number of camera views, SV4D outputs a $V{\mkern-2.0mu\times\mkern-2.0mu}F$ grid of images, as illustrated in Fig. 1. Given a reference video, we first obtain the reference multi-views for the first frame using an off-the-shelf multi-view generator (SV3D (Voleti et al., 2024)). SV4D then jointly outputs the remaining grid of images as highlighted by a yellow box in Fig. 1. In contrast to prior and concurrent works, SV4D jointly reasons along both view and motion axes, resulting in state-of-the-art multi-frame and multi-view consistency in the output novel view videos.
Specifically, we start with Stable Video Diffusion (SVD) (Blattmann et al., 2023a), a state-of-the-art video generator, and equip it with two attention blocks: view attention and frame attention. The view attention block aligns the multi-view images at each video frame, conditioning on the first view i.e, in the reference video. Similarly, the frame attention block aligns the multi-frame images at each view, conditioning on the first frame at each view, i.e, reference multi-view. This dual-attention design leads to significantly improved dynamic and multi-view consistency compared to prior art.
@@
Given the lack of large-scale 4D datasets, we carefully initialized SV4D weights with those from SVD (Blattmann et al., 2023a) and SV3D (Voleti et al., 2024) networks, thereby leveraging the priors learned in the existing video and multi-view diffusion models. To further train SV4D, we carefully curated a subject of Objaverse (Deitke et al., 2023b; a) dataset with dynamic 3D objects, resulting in the ObjaverseDy dataset.
We perform extensive comparisons of both novel view video synthesis and 4D generation results with respective state-of-the-art methods on datasets with synthetic (ObjaverseDy, Consistent4D (Jiang et al., 2024c)), and real-world (DAVIS (Perazzi et al., 2016; Pont-Tuset et al., 2017; Caelles et al., 2019)) data. We modify the FVD metric (Unterthiner et al., 2018) to evaluate both video frame and view consistency, validating the effectiveness and design choices of our approach. Fig. 1 shows some sample results of our approach.
@@ 2 Related Work
3D Generation. Here, we refer to the works that generate static 3D content as 3D generation. DreamFusion (Poole et al., 2023) first proposed to distill priors from the 2D diffusion model via SDS loss to optimize the 3D content from text or image. Several subsequent works (Yi et al., 2024; Tang et al., 2024; Shi et al., 2024a; Wang et al., 2024b; Li et al., 2024b; Weng et al., 2023; Pan et al., 2024a; Chen et al., 2024b; Sun et al., 2024; Sargent et al., 2024; Liang et al., 2024b; Zhou et al., 2024; Guo et al., 2023) try to solve issues caused by the original SDS loss, such as multi-face Janus, slow generation speed, and over-saturated/smoothed generations. Recent works (Hong et al., 2024; Jiang et al., 2024a; Wang et al., 2024a; Zou et al., 2024; Wei et al., 2024; Tochilkin et al., 2024) try to directly predict the 3D model of an object via a large reconstruction model. Another approach (Liu et al., 2023b; 2024b; Long et al., 2024; Voleti et al., 2024; Ye et al., 2024; Karnewar et al., 2023; Li et al., 2024a; Shi et al., 2024b; 2023; Wang & Shi, 2023; Liu et al., 2023a; 2024a) to 3D generation is generating dense multi-view images with sufficient 3D consistency. 3D content is reconstructed based on the dense multi-view images. We follow this strategy, but generate consistent multi-view videos (instead of images) and then reconstruct the 4D object.
@@
Video Generation. Recent video generation models (Ho et al., 2022; Voleti et al., 2022; Blattmann et al., 2023b; a; He et al., 2022; Singer et al., 2023a; Guo et al., 2024) have shown very impressive performance with consistent geometry and realistic motions. Video generation models have good generalization capabilities, as they are trained on large-scale image and video data that are easier to collect than large-scale 3D or 4D data. Hence, they are commonly used as foundation models for various generation tasks. Well-trained video generative models have shown their potential to generate multi-view images as a 3D generator (Voleti et al., 2024; Chen et al., 2024c; Han et al., 2024; Kwak et al., 2024; Melas-Kyriazi et al., 2024). SV3D (Voleti et al., 2024) adapts SVD (Blattmann et al., 2023a) to generate novel multiple views. In this work, we leverage the pre-trained video generation model for 4D generation by adding an additional view attention layer to align the multiview images.
@@
4D Generation. On one hand, recent optimization-based methods (Singer et al., 2023b; Bahmani et al., 2024; Ling et al., 2024; Ren et al., 2023; Zhao et al., 2023; Jiang et al., 2024c; Zeng et al., 2024; Yin et al., 2023; Chen et al., 2024a) can generate 4D content by distilling pre-trained diffusion models in a 4D representation (Cao & Johnson, 2023; Kerbl et al., 2023; Mildenhall et al., 2021) via SDS loss Poole et al. (2023). However, they tend to take hours to generate 4D content due to time-consuming optimization. On the other hand, photogrammetry-based methods (Yang et al., 2025; Pan et al., 2024b; Sun et al., 2025) mimic a 3D object capture pipeline by directly generating multi-frame multi-view images of a 4D content with dynamic and multi-view consistency and then directly reconstructing 4D representations with them. Although these methods have much faster speeds, inference-only pipelines are adopted due to the paucity of the 4D data, thus making the spatial-temporal consistency still unsatisfactory. Moreover, some models like EG4D (Sun et al., 2025) further rely on a third model (SDXL-Turbo (Sauer et al., 2025)) to refine the 4D assets. In this work, we proposed a monolithic model to generate a more consistent image grid, and we used a carefully curated 4D dataset to train the model. Recently, there has been a surge of interest in 4D generation, resulting in the development of concurrent works (Liang et al., 2024a; Zhang et al., 2024; Li et al., 2025; Ren et al., 2025; Wang et al., 2025; Jiang et al., 2024b). Diffusion4D (Liang et al., 2024a) only generates diagonal images (space-time), by fine-tuning a 4D-aware video diffusion model with motion magnitude guidance. Vivid-ZOO (Li et al., 2025) combines the 3D and video generative model and then finetunes alignment modules to mitigate the incompatibility between reused layers. 4Diffusion (Zhang et al., 2024) inserts temporal layers to a 3D-aware diffusion model to generate novel view videos. Both Vivid-ZOO and 4Diffusion freeze pre-trained weights, limiting their ability to fully capture spatial-temporal information. Additionally, they generate low-resolution images or a limited number of frames due to memory constraints. In contrast, SV4D finetunes all layers of our unified model, allowing for comprehensive spatial-temporal learning, and uses a mixed sampling scheme, enabling the generation of high-resolution and long novel views.
@@ 3 Method
Our main idea is to build multi-frame and multi-view consistency in a 4D object by surgically combining the frame-consistency in a video diffusion model, with the multi-view consistency in a multi-view diffusion model. In our case, we choose SVD (Blattmann et al., 2023a) and SV3D (Voleti et al., 2024) as the video and multi-view diffusion models respectively, for the advantages reasoned below. However, it is to be noted that any choice of attention-based diffusion models should work.
@@ 3.1 Novel View Video Synthesis via SV4D
Problem Setting. Formally, we begin with a monocular input video ${\bf J}\in\mathbb{R}^{F{\mkern-2.0mu\times\mkern-2.0mu}D}$ of a dynamic object of $F$ frames and $D\coloneq 3{\mkern-2.0mu\times\mkern-2.0mu}H{\mkern-2.0mu\times\mkern-2.0mu}W$ dimensions, where $H$ and $W$ are the height and width of each frame. Our goal is to generate an image matrix ${\bf M}\in\mathbb{R}^{V{\mkern-2.0mu\times\mkern-2.0mu}F{\mkern-2.0mu\times\mkern-2.0mu}D}$ of the 4D object consisting of $V$ camera views and $F$ dynamic frames of each frame. Similar to SV3D (Voleti et al., 2024), the multi-view frames follow a camera pose trajectory $\bm{\pi}\in\mathbb{R}^{V{\mkern-2.0mu\times\mkern-2.0mu}2}=\{(e_{v},a_{v})\}_{v=1}^{V}$ as a sequence of $V$ tuples of elevation $e$ and azimuth angles $a$. We generate this image matrix by iteratively denoising samples from a learned conditional distribution $p({\bf M}|{\bf J},\bm{\pi})$, parameterized by a 4D diffusion model. Due to memory limitation in generating $V{\mkern-2.0mu\times\mkern-2.0mu}F$ images simultaneously, we break down the full generation process into multiple submatrix generation steps.
@@
SV4D Network. Our goal is to make the generated image matrix $\bf M$ dynamically consistent in the “frame” axis, and multi-view consistent in the “view” axis (see Fig. 2 left). To achieve this, we condition the image matrix generation on the corresponding frames of the monocular input video $\{{\bf M}_{0,f}\}={\bf J}$ as well as the views from reference multi-view images of the first video frame $\{{\bf M}_{v,0}\}$, which provide motion and multi-view information of the 4D object, respectively. Without loss of generality, we obtain the reference multi-view images of the first input frame by sampling from a pre-trained SV3D (Voleti et al., 2024) model, which can be expressed as $p(\{{\bf M}_{v,0}\}|\{{\bf M}_{0,0}\},\bm{\pi})$. The overall SV4D novel view video synthesis can be rewritten as sampling from the distribution $p({\bf M}|\{{\bf M}_{0,f}\},\{{\bf M}_{v,0}\},\bm{\pi})$.
@@
Figure 2: SV4D model architecture. For camera conditioning, we feed the sinusoidal embedding of camera viewpoints to the convolutional blocks in the UNet, and use the input video for cross-attention conditioning in the spatial and view attention blocks. To improve temporal consistency, we introduce an additional frame attention block, conditioned on the corresponding views of the first frame.
@@
We build the SV4D network based on SVD (Blattmann et al., 2023a) and SV3D (Voleti et al., 2024) models to combine the advantages of both video and multi-view diffusion models. As illustrated in Fig. 2, SV4D consists of a UNet with multiple layers, where each layer contains a sequence of one residual block with Conv3D layers and three transformer blocks with attention layers: spatial, view, and frame attention. Similar to SV3D, the residual Conv block takes in the noisy latents of flattened image matrix as well as handles the incorporation of conditioning camera poses $\{\bm{\pi}_{v}\}$, and the spatial attention layer handles image-level details by performing attention across the image width and height axes. To better capture motion cues in the input monocular video $\{{\bf M}_{0,f}\}$, we concatenate its VAE latents to the noisy latents ${\bf z}_{n}$ before feeding it to the UNet.
@@
To further ensure dynamic consistency across video frames, we insert a frame attention layer in each UNet block, which applies the attention mechanism in the video frame dimension. The frame attention of each novel view video is conditioned on the corresponding reference view (of the first frame) $\{{\bf M}_{v,0}\}$ via cross-attention, allowing the network to preserve dynamic coherence starting from the first frame. We initialize the weights of the frame attention layers from SVD and the rest of the network from SV3Dp, to leverage the generalizability as well as rich dynamic and multi-view priors learned from large-scale video and 3D datasets. More network details are in Appendix A.4.
ObjaverseDy Dataset. Considering that there exists no large-scale training datasets with dynamic 3D objects, we curate a new 4D dataset from the existing Objeverse dataset (Deitke et al., 2023b; a), a massive dataset with annotated 3D objects. Objaverse includes animated 3D objects, however, several of these animated 3D objects are not suitable for training due to having too few animated frames or insufficient motion. In addition, in the rendering stage, the common rendering and sampling setting may cause some issues. For example, dynamic objects may be out of the image if the camera distance is fixed because they have global motion; the motion of objects may be too fast or too slow if the temporal sampling step is fixed.
We follow several steps to curate and clean the 4D objects for our training purposes. We first filter out the objects based on a review of licenses. Then, we remove the objects whose animated frames are too few. To further filter out objects with minimal motion, we subsample keyframes from each video and apply simple thresholding on the maximum L1 distance between these frames as motion measurement. To render the training novel view videos, we flexibly choose the camera distance from the object. Starting from a base value, we increase the camera distance until the object fits within all frames of the rendered images. We also dynamically adjust the temporal sampling step. Starting from a base value, we increase the sampling step until the L1 distance between consecutive keyframes exceeds a certain threshold. These steps ensure a high-quality collection of 4D objects, with rendered multi-view videos that form our ObjaverseDy dataset. More dataset details are in Appendix A.1.
Training Details. We train SV4D on our ObjaverseDy dataset. We choose to finetune from the SV3Dp (Voleti et al., 2024) model to output 40 frames ($F=5$ frames along each of the $V=8$ views) with the spatial resolution $576\times 576$, where the parameters in the frame attention layers are initialized from SVD-xt (Blattmann et al., 2023a). Similar to SV3D, we train SV4D progressively by first training on the static camera orbits for 40K iterations, then fine-tuning it for 20K iterations on the dynamic orbits. We use an effective batch size of 16 during training on 2 nodes of 8 80GB H100 GPUs. For more training details, please see Appendix A.2.
@@
Inference Sampling Scheme. Due to memory limitations, we cannot generate all the novel view frames at once. To generate the full $V{\mkern-2.0mu\times\mkern-2.0mu}F$ image matrix with arbitrary length videos, one can naively run the submatrix generation independently. However, we observe that it often leads to severe artifacts due to the inconsistencies between consecutive submatrices. Hence, we design a novel sampling scheme to mitigate the issue. The proposed sampling scheme is illustrated in Fig. 3. Note that we only show motion frame extensions here for illustrative purposes. We first generate a sparse set of anchor frames with SV4D (interleaved sampling) as shown in Fig. 3 (left). Then we use the anchor frames as new reference views to densely sample the remaining frames (dense sampling). To ensure a smooth transition between consecutive generations, we alternatively use the first (forward) or last (backward) anchor frame for conditioning at each diffusion step, as shown in Fig. 3 (middle). In experiments (Fig. 8), we demonstrate that our sampling scheme can generate novel view videos that are more temporally consistent compared to using an off-the-shelf video interpolation model (Li et al., 2023) to interpolate in between the SV4D-generated anchor frames.
@@
4D Representation. As shown in Fig. 4, given the novel view videos generated by SV4D, we optimize a 4D representation to reconstruct the dynamic 3D asset. Formally, we learn a neural representation $\Psi_{\theta}:(\mathbf{x},t)\mapsto(\sigma,c)$ that maps the sampled 3D points $\mathbf{x}=(x,y,z)$ along a camera ray, and the time embedding $t$ (i.e. continuous version of “frame” in the earlier discrete case) to its volumetric density $\sigma\in\mathbb{R}_{+}$ and color $c\in\mathbb{R}^{3}_{+}$. Similar to D-NeRF (Pumarola et al., 2021), we represent a 4D object by the composition of a canonical NeRF that captures the static 3D appearance and a deformation field which handles the object motion across time. For each 3D point $\mathbf{x}$ in the canonical space, we trilinearly interpolate the multi-resolution hash grid features following Instant-NGP (Müller et al., 2022), and decode them as density and color via MLPσ and MLPc, respectively. The deformation field is represented by an MLP network MLPt conditioned on time embedding $t$, mapping temporal deformation of $\mathbf{x}$ to the common canonical space $\mathbf{x}+\Delta\mathbf{x}$. Overall, the dynamic NeRF parameters $\theta$ include the canonical hash grid, MLPt, MLPσ, and MLPc. Since we do not exhaustively sample temporal timestamps or dense spatial views like in prior SDS-based approaches, we observe that this dynamic NeRF representation produces better 4D results compared to other representations such as 4D Gaussian Splatting (Wu et al., 2024), which suffers from flickering artifacts and does not interpolate well across time or views. We show several examples to demonstrate this in the supplementary rebuttal video.
Optimization Details. By leveraging the consistent image matrix generated by SV4D as pseudo ground-truths, we adopt a simple photometry-based optimization without the cumbersome SDS losses. The reconstruction losses include a pixel-level MSE loss, mask loss, and a perceptual LPIPS (Zhang et al., 2018b) loss. Similar to SV3D (Voleti et al., 2024), we also use several geometric priors to regularize the output shapes, such as a mono normal loss similar to MonoSDF (Yu et al., 2022) as well as bilateral depth and normal smoothness losses (Voleti et al., 2024) to encourage smooth 3D surfaces where the projected image gradients are low.